Mr. Newman needs some repair work done on his truck. He gets two estimates for the repairs. Itza-Reck Autos charges $162 for parts and $32 per hour for labor. Cars-R-Us will charge $144 for parts and $35 per hour for labor. Which is the better deal?
Answer by checkley71(8403)

TO SOLVE THIS PROBLEM WE NEED TO KNOW HOW MANY HOURS ARE REQUIRED FOR THE REPAIR WORK FROM EACH ESTIMATOR.
162+32X=? WHERE X IS THE NUMBER OF HOURS FOR THE REPAIRS.
144+35X=?
----------------------------------------------
IF X=1 THEN 162+32=194 & 144+35=179 ITZA BEST.
IF X=5 THEN 162+32*5=162+160=322 & 144+35*5=144+175=319 ITZA BEST.
IF X=8 THEN 162+32*8=162+256=418 & 144+35*8=144+280=424 CARS BEST.